OpenStudy (anonymous):

Let f(x) = x2 + 3x - 4 and g(x) = x + 5. Find f(x) · g(x). x3 + 3x2 + 16x - 20 x3 + 5x2 + 14x - 20 x3 + 8x2 + 11x - 20 x3 + 9x2 + 19x - 20

OpenStudy (anonymous):

could you use FOIL on this?

OpenStudy (anonymous):

also x2 is x^2

OpenStudy (anonymous):

ok so if f= x^2+3x-4 and g=x+5 and we need to find f*g we must set it up like :

OpenStudy (anonymous):

do you know where to go from here?

OpenStudy (anonymous):

I'm lost lol

OpenStudy (anonymous):

I see it

OpenStudy (anonymous):

You ever heard of a punnet square? Well that is what it is, You'll get the same answer as you would using foil, but I understand this way better so I used it like this. DO you understand how to see foil then?

OpenStudy (anonymous):

Yeah I see FOIL now

OpenStudy (anonymous):

so now do we combine like terms?

OpenStudy (anonymous):

yes, :) and you'll see your answer up there